### v1

# Reversibility in finite-dimensional collapse dynamics

Abstract
We study finite-dimensional quantum systems with arbitrary, discontinuous collapse events. Formally, we fix a realization map (a physically admissible selector of the collapse dynamics) and do not rely on any regularity of the induced dynamics. We prove the existence of a topologically closed, invariant subset of the projective state space in which any two states can be connected with arbitrarily fine Fubini-Study precision and arbitrarily small integrated energetic cost. This shows that the preservation of information along a realized branch guarantees islands of quasi-reversibility, while genuine irreversibility requires additional ingredients such as non-compactness, explicit erasure, or coupling to reservoirs.
KEYWORDS: Quantum collapse dynamics; Quasi-reversibility; Chain-recurrence; Information non-erasure.
MSC2020: 81P15; 81P45; 37B20; 03E10.
1 Introduction
1.1 Reversibility in unitary versus collapse dynamics
In unitary quantum dynamics of finite-dimensional systems, recurrence phenomena, either in the measurable or topological sense, arise naturally from the continuity of unitary evolution. Indeed, since the Schrödinger dynamics is continuous (indeed isometric), in the compact case the classical recurrence theorems by Poincaré [poincare1890probleme, bocchieri1957quantum] and Birkhoff [birkoff1912quelques] ensure, respectively, that almost every orbit admits arbitrarily close returns in state space, and that quasi–periodic orbits exist. Therefore, within purely unitary evolution of compact systems, the dynamics admits no intrinsic arrow of time.
When wavefunction collapse is included in the picture, this structure breaks. For the discrete–time setting, a natural analogue of the previously described unitary evolution is obtained by following a single, consistent branch of the dynamics in an open system, an approach that has been introduced for the first time (within quantum optics) in [dalibard1992wave], and subsequently elaborated in many other works (for instance [carmichael1993open] and [wiseman2009quantum], the latter also discussing single-branch dynamics from an epistemological point of view). Consider a finite-dimensional space $\mathbb{P}(\mathcal{H})$ of pure states, a single observable $A$ with finite spectrum, and a particular state $\psi$. At each time step either no collapse occurs (so one applies a unitary Schrödinger evolution $\psi\mapsto U\psi$) or a projective collapse $\{P_{i}\}$ occurs. The future outcome itinerary that is realized by $\psi$, that we indicate by $\omega=\{i_{0},i_{1},\dots\}$ ($i_{k}$ being symbolic tools representing the possible collapse outcomes as well as the “blank”, no-collapse case), has to be of course admissible from a physical point of view. In particular, each selected outcome must have a nonzero Born weight for the state immediately preceding that step. If we assign consistently a selected outcome to every state, the evolution along the outcome trajectory provides a map $T:\mathbb{P}(\mathcal{H})\to\mathbb{P}(\mathcal{H})$ that is, in general, many–to–one and highly discontinuous, because there is no justification in assuming that close states will evolve into close collapsed states. Arbitrarily close states are allowed, in general, to collapse to eigenstates belonging to distinct eigenspaces of the chosen observable; this is not at all surprising if one considers that even the same pre–collapse state may be sent to different post–collapse eigenstates at distinct collapse events. Moreover, the eigenspaces of the observable are not absorbing for $T$ unless $U$ preserves them (i.e. $U,P_{i}]=0$); between collapses, the unitary evolution typically carries states out of these subspaces. Consequently, the hypotheses behind the Poincaré/Birkhoff recurrence theorems are violated on the observed system. [^2] If one re–includes apparatus and environment as a closed finite-dimensional composite, the global pure state again evolves unitarily on a compact projective space, and recurrence can be recovered.
In the classical compact setting (where the Poincaré–Birkhoff recurrence theorems apply), one studies the long–time behavior of a single map on a compact state space. We want to explore the closest possible analogue when generalizing unitary evolution of closed systems to collapse dynamics of open systems. We fix a selector $\omega$ of the skew-product dynamics and thus obtain a single branch on the projective state space. Notice that this choice cleanly separates the predictability issue (which outcome occurs) from the regularity issue (whether the induced map $T$ is continuous). Moreover, we take the selector in its most literal observational sense:
  1. 1.
    every outcome itinerary has to be physically admissible and time–compatible for each initial state;
  2. 2.
    each selected outcome has to have positive Born weight at that step;
  3. 3.
    finally, treating collapse as intrinsically unpredictable at the single-event level [^3] As it is assumed in the standard von Neumann–Dirac formulation. Alternative interpretations, such as Bohmian mechanics or the Everett (many-worlds) interpretation, eliminate collapse entirely; the present work follows the conventional projective measurement framework., there is no basis to assume that an entire open neighborhood is shielded from any outcome with nonzero Born weight, which means that we have, in general, topological accessibility of every outcome label arbitrarily close to every state.
We remark that is not the standard stance in the quantum–trajectories literature: the seemingly intractable irregularity, that such state–wise selectors permit, is usually avoided by averaging over outcomes or retaining stochasticity and works with the skew product or with stochastic master equations (diffusive/jump SMEs), where continuity and stability are recovered in measure or in expectation rather than pointwise for a deterministic selector–induced map [carmichael1993open, belavkin1992quantum, barchielli2009quantum, attal2006repeated, wiseman2009quantum]. Our approach is instead to explore what can be said, from the point of view of recurrence, even allowing full irregularity. We fix thus a realization map (the chosen selector of the dynamics) assuming only the (physics-dictated) properties 1-2 above, and allowing the density property (3), which implies everywhere discontinuity for the induced evolution map $T$.
1.2 The result
Once a realization map $D$ has been chosen, and an ambient Hamiltonian $H$ is fixed, with corresponding unitary group $U$, they determine a (very irregular) induced dynamics
$T_{U,D}:\mathbb{P}(\mathcal{H})\longrightarrow\mathbb{P}(\mathcal{H}),$
where points are unit vectors $u]\in\mathbb{P}(\mathcal{H})$.
Landauer’s principle (see [landauer1961irreversibility]) bounds the minimal heat cost of logically irreversible operations (such as erasure), and conversely permits, in principle, arbitrarily low dissipation for logically reversible transformations in the quasi–static limit.
The choice of a selector for the dynamics can be interpreted as realizing the extreme, no-information-erasure limit of Landauer’s principle, in the sense highlighted by C.H. Bennett’s resolution of the Maxwell demon paradox ([bennett1982thermodynamics, bennett2003notes, hanson2018landauer]). In particular, we share the following position, expressed by Bennett (from [bennett1982thermodynamics]):
“[…] the essential irreversible step, which prevents the demon from breaking the second law, is not the making of a measurement (which in principle can be done reversibly) but rather the logically irreversible act of erasing the record of one measurement to make room for the next.”
Our realization map, indeed, does not erase (or coarse-grain) outcome information: the entire future outcome sequence is retained as a fixed record, and collapse consists solely in following one realized branch.
As a physical principle, however, Landauer’s is permissive rather than generative: it does not assert that a given dynamics actually realizes quasi–reversibility, nor does it identify when it must occur. In the present work we show that, despite the absence of regularity assumptions on $T$, there exists a topologically closed, TT–invariant, strongly chain–transitive subset of projective state space along which any two states can be connected with arbitrarily small Fubini–Study error and arbitrarily small integrated energetic cost (in the sense of Def. 4.1). While Landauer’s principle says that, without erasure, quasi–reversibility is not thermodynamically forbidden, we prove that in finite dimension, an island of operational reversibility is topologically enforced under no erasure.
Let us make this more precise. For each finite forward itinerary
$\gamma=\big([u_{0}],[u_{1}],\dots,[u_{n}]\big),$
we denote by
$\mathcal{E}_{D}(\gamma)\,\geq 0$
its energetic cost, that is the energy needed to modify the ambient Hamiltonian in order to let the state $[u]$ evolve along $\gamma$ (this will be made precise in Def. 4.2).
Given two states $u],[v]\in\mathbb{P}(\mathcal{H})$, we can define the minimal forward cost of going from $[u]$ to $[v]$ under the dynamics law $T_{U,D}$ by
[…]
Definition 1.1.
We say that the dynamics $(T_{U,D},\mathcal{E}_{D})$ exhibits an operational arrow of time if there exist states $u],[v]\in\mathbb{P}(\mathcal{H})$ such that
$d_{D}\big([u]\to[v]\big)\;<\;d_{D}\big([v]\to[u]\big),$
i.e. the energetic effort required by the microscopic law to carry the system from $[u]$ to $[v]$ in forward time is strictly smaller than the minimal energetic effort required to realize the reverse change from $[v]$ back to $[u]$.
Conversely, on a subset $B\subset\mathbb{P}(\mathcal{H})$ we say that the operational arrow of time does not arise (or that $B$ is operationally reversible) if
$d_{D}\big([u]\to[v]\big)\;=\;d_{D}\big([v]\to[u]\big)\;=\;0\qquad\text{for all }[u],[v]\in B.$
We can summarize our main result as follows:
Main result.
In the no–erasure regime, compactness enforces quasi–reversibility: there exists a recurrent, invariant subsystem $S$ allowing quasi-returns with arbitrarily small error and arbitrarily small energetic cost.
In particular, collapse alone does not generate an operational arrow of time on $S$.
Genuine irreversibility, thus, requires additional ingredients: information loss/coarse–graining, non-compact limits, or coupling to reservoirs.
1.3 Where a naive diagonal grid argument fails
It may look tempting to try to get a (naive) heuristic argument proving our main result as follows: fix $\epsilon>0$, cover $\mathbb{P}(\mathcal{H})$ by finitely many $d_{FS}$–balls ($\epsilon$–cells”), and note that along any long enough forward segment some cell must be visited twice; this yields an $\epsilon$–loop. Pushing $\epsilon\to 0$ seems to promise finer and finer quasi-loops. However, this argument is intrinsically scale–wise: as $\epsilon$ shrinks, the repeating cell and the base point inside it drift with $\epsilon$. Of course, by compactness the chosen cells admit a convergent subsequence, but in the absence of continuity the induced dynamics at the limit point is decoupled from the finite $\epsilon$–loops: $T$ at the limit may bear no relation to the images seen along the approximants. Consequently, the grid argument does not deliver a chain–recurrent point, let alone a topologically closed, invariant, internally strongly chain–transitive subsystem; nothing in this argument enforces nesting across scales of the pseudo–orbits.
To make this remark more precise, fix a mesh $\epsilon_{k}=2^{-k}$ and choose refining covers $\mathcal{U}_{k+1}\prec\mathcal{U}_{k}$ by $d_{FS}$–balls of radius $\epsilon_{k}$. At scale $k$, scan a long forward segment and pick the first cell $C_{k}\in\mathcal{U}_{k}$ that is visited twice; let $\Gamma_{k}$ be the loop between those visits. This yields strong $\epsilon_{k}$–loops, but the selected cells need not nest: typically $C_{k+1}\nsubseteq C_{k}$. Even after extracting a nested subsequence $C_{k_{r}}\searrow\{x_{\ast}\}$ (via a refinement tree), the discontinuity of $T$ decouples the local dynamics at $x_{\ast}$ from the images traced by the $\Gamma_{k_{r}}$, so neither $x_{\ast}\,\mathcal{SC}_{d_{FS}}\,x_{\ast}$ (where $\mathcal{SC}_{d_{FS}}$ is the strong chain-relation defined in Def. 2.3) nor an invariant, internally $\mathcal{SC}_{d_{FS}}$–transitive set is guaranteed. Summarizing, the grid-refining technique only shows that:
For every $\epsilon$, there is a cell revisited with $O(\epsilon)$ accuracy and energetic cost,
while we search for the stronger property:
There is a point such that, for every $\epsilon$, it is revisited with $O(\epsilon)$ accuracy and energetic cost.
We implement (in Appendix) a transfinite argument that enforces both spatial nesting and dynamical coherence across scales: at successor stages it follows the dynamics induced by the selector; at limit stages it passes to accumulation points of cofinal sequences, gluing thus information from all previous scales. This procedure yields a strongly chain–recurrent point, and even a topologically closed TT–invariant subset that is internally $\mathcal{SC}_{d_{FS}}$–transitive. We remark that the procedure is constructive: we do not use any form of the Axiom of Choice and the two basic operations considered (taking images of the map and passing to an accumulation point of an orbit) clearly have finite-precision analogues.
2 Preliminaries
2.1 Notation and definitions for Quantum Dynamics formalism
Let $\mathcal{H}\cong\mathbb{C}^{n+1}$ be a finite-dimensional Hilbert space (with $n\geq 0$). Note that $\dim_{\mathbb{C}}\mathcal{H}=n+1$, so the associated complex projective space $\mathbb{P}(\mathcal{H})\cong\mathbb{CP}^{n}$ has complex dimension $n$. Let $\mathbb{P}(\mathcal{H})$ be the complex projective space of one-dimensional subspaces (pure states), equipped with the Fubini–Study metric $d_{FS}$ (see Def. 2.1 below), which makes $\mathbb{P}(\mathcal{H})$ a compact metric space.
We recall the definition of the Fubini-Study metric (for a more detailed discussion, see [bengtsson2017geometry]).
Definition 2.1 (Fubini–Study metric).
Let $\mathbb{CP}^{n}$ be the complex projective space of complex dimension $n$ (note that $\mathbb{P}(\mathbb{C}^{n+1})$ is an equally standard notation). A point in $\mathbb{CP}^{n}$ can be represented by homogeneous coordinates $z_{0}:z_{1}:\ldots:z_{n}]$, where $(z_{0},z_{1},\ldots,z_{n})\in\mathbb{C}^{n+1}\setminus\{0\}$ and
$z_{0}:z_{1}:\ldots:z_{n}]=[\lambda z_{0}:\lambda z_{1}:\ldots:\lambda z_{n}]\quad\text{for all }\lambda\in\mathbb{C}\setminus\{0\}.$
[…]
representing the points $\phi],[\psi]\in\mathbb{CP}^{n}$, respectively. Then the Fubini–Study distance between $\phi]$ and $\psi]$ is defined by
[…]
Equivalently, the Fubini–Study metric is the Kähler metric whose associated $(1,1)$-form is
$\omega_{FS}=\frac{i}{2}\,\partial\bar{\partial}\log\|z\|^{2},\qquad\text{where }\|z\|^{2}=\sum_{j=0}^{n}|z_{j}|^{2}.$
Denote by $\mathrm{U}(\mathcal{H})$ the set of all linear unitary operators from $\mathcal{H}$ to itself. Fix $U\in\mathrm{U}(\mathcal{H})$ and a finite collection of orthogonal projections $\{P_{j}\}_{j\in\Lambda}$ on $\mathcal{H}$, indexed by a finite alphabet
$\Lambda=\{0,1,\dots,m\}.$
We do not assume any property for the projections $P_{j}$ beyond, linearity, orthogonality and idempotence. We interpret these projections $\{P_{j}\}_{j\in\Lambda}$ as the spectral projections of a single self-adjoint observable $A$ on $\mathcal{H}$: there exist distinct eigenvalues $\{\lambda_{j}\}_{j=1}^{m}$ such that
$A=\sum_{j=1}^{m}\lambda_{j}P_{j},\qquad P_{j}P_{k}=\delta_{jk}P_{j},\qquad\sum_{j=1}^{m}P_{j}=I,$
where $P_{j}$ projects onto the (possibly degenerate) eigenspace $E_{\lambda_{j}}=\operatorname{im}P_{j}$. The index 0 is reserved for the “blank” (no-collapse) channel, and therefore we set $P_{0}:=I$; it is not part of the spectral resolution of $A$.
The unitary operator $U$ represents the “ambient” free continuous evolution between collapse events. Let $U(t_{1},t_{0})$ denote the unitary evolution through $U$ from time $t_{0}$ to time $t_{1}$. If the Hamiltonian $H$ is time-independent, then
$U(t_{1},t_{0})=e^{-iH(t_{1}-t_{0})},$
where we use the normal units: $\hbar=1.$ If $H$ depends on time, i.e. $H=H(t)$, the evolution is given by the time-ordered exponential
$U(t_{1},t_{0})=\mathcal{T}\exp\left(-i\!\int_{t_{0}}^{t_{1}}H(s)\,ds\right),$
which satisfies
$U(t_{2},t_{1})U(t_{1},t_{0})=U(t_{2},t_{0})\text{ for $t_{0}<t_{1}<t_{2},$}$
$U(t_{1},t_{0})^{\dagger}U(t_{1},t_{0})=I,$
$U(t,t)=I,$
where $I$ is the identity operator. In the following, when we have $t_{0}=0$, we use the notations $U(t)\coloneqq U(t,0)$ and also set $U\coloneqq U(1)=U(1,0)$.
A single collapse event with outcome index $j\in\Lambda$ acts (where defined) by
$u]\mapsto f_{j}([u])\coloneqq\frac{P_{j}Uu}{\lVert P_{j}Uu\rVert},$
where $u\in\mathcal{H}$ is any unit vector that represents the point $u]\in\mathbb{P}(\mathcal{H})$, and the expression is defined when $\lVert P_{j}Uu\rVert\neq 0$.
2.2 Notation and definitions for abstract Topological Dynamics formalism
In this Section we provide the topological dynamical concepts that will be used to prove our main result. Most of them appear only in the proof of Theorem 4.5, given in the Appendix.
We say that $(X,f)$ is a topological dynamical system if $X$ is a topological space with $d:X^{2}\to\mathbb{R}_{0}^{+}$ a metric compatible with the topology on $X$ and $f:X\to X$ a map. By $f^{n}(x)$ with $x\in X$ and $n\in\mathbb{N}$ we mean the $n$-times composition $f\circ f\circ\ldots\circ f(x)$. When we say that $(X,f)$ is a compact dynamical system we require, in addition, that $X$ is a compact metric space.
Let us recall the notion of $\epsilon$-chain or $\epsilon$-pseudo-orbit [kurka2003topological, p. 48].
Definition 2.2.
Given two points $x,y\in X$ and $\epsilon>0$, an $\epsilon$-chain (also called $\epsilon$-pseudo-orbit) from $x$ to $y$ is a finite set of points $x_{0},x_{1},\ldots,x_{n}$ in $X$, with $n\geq 1$, such that
  1. i)
    $x_{0}=x$ and $x_{n}=y$,
  2. ii)
    $d(f(x_{i}),x_{i+1})<\epsilon$ for every $i=0,1,\ldots n-1$.
The chain relation $\mathcal{C}\subseteq X^{2}$ is the binary relation defined as follows: given $x,y\in X$,
[…]
Let $\mathcal{A}\subseteq X^{2}$ be a binary relation on $X$. For $N\subseteq X$ and $y\in X$, we write $N\,\mathcal{A}\,y$ if $x\,\mathcal{A}\,y$ for every $x\in N$. Let $CR(X,f)=\{x\in X:x\,\mathcal{C}\,x\}$ be the set of all the chain-recurrent points of $(X,f)$ (we write simply $CR_{f}$ when the space $X$ is clear from the context).
Recall the concept of strong chain-recurrence originally introduced by Easton [easton2006chain].
Definition 2.3.
Given two points $x,y\in X$ and $\epsilon>0$, we say that a finite sequence of points $x_{0},x_{1},\ldots,x_{n}$ of $X$, with $n\in\mathbb{N}$, is a strong $(\epsilon,d)$-chain from $x$ to $y$ if
  1. i)
    $x_{0}=x$ and $x_{n}=y$,
  2. ii)
    $\sum_{i=0}^{n-1}d(f(x_{i}),x_{i+1})<\epsilon$.
The strong chain relation $\mathcal{SC}_{d}\subseteq X^{2}$ is the binary relation defined as follows: given $x,y\in X$,
[…]
Let $\mathcal{SCR}_{d}(f)=\{x\in X:x\,\mathcal{SC}_{d}\,x\}$ be the set of all the strong chain-recurrent points of $(X,f)$ (we write simply $\mathcal{SCR}_{d}$ when the map $f$ is clear from the context).
Definition 2.4.
We say that $S\subseteq X$ is $\mathcal{SC}_{d}$-transitive if for every $x,y\in S$ we have $x\,\mathcal{SC}_{d}\,y$. If, in addition, for every $\epsilon>0$, there exists a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $S$ we say that $S$ is internally $\mathcal{SC}_{d}$-transitive.
Definition 2.5.
The derived set $S^{\prime}$ of a subset $S\subseteq X$ is the set of the limit points $x$ of $S$, that is, the set of points $x\in X$ such that for every neighborhood $U$ of $x$ we have $S\cap(U\setminus\{x\})\neq\emptyset$.
Throughout the paper we use the standard Landau notation $O(\epsilon^{k})$ to denote terms whose norm is bounded by $C\,|\epsilon|^{k}$ for some constant $C>0$ independent of $\epsilon$.
3 Discrete-time Quantum Dynamics
Definition 3.1.
For each $j\in\Lambda$, set
$D_{j}:=\{[u]\in\mathbb{P}(\mathcal{H}):\ \lVert P_{j}Uu\rVert\neq 0\}.$
Equivalently, $D_{j}$ is the set of pure states for which the outcome $j$ has strictly positive Born probability after applying $U$. Notice that $D_{0}=\mathbb{P}(\mathcal{H})$.
For each $j\in\Lambda$ the map $f_{j}:D_{j}\to\mathbb{P}(\mathcal{H})$ defined by $f_{j}([u])=\dfrac{P_{j}Uu}{\lVert P_{j}Uu\rVert}$ is continuous on $D_{j}$.
Let $\Omega=\Lambda^{\mathbb{N}_{0}}=\{(w_{n})_{n\in\mathbb{N}_{0}}:w_{n}\in\Lambda\text{ for all }n\in\mathbb{N}_{0}\}$ be the one-sided full shift with the product topology (equivalently the metric $d_{\Omega}(\omega,\omega^{\prime})=2^{-N}$ where $N$ is the largest positive integer such that $\omega_{n}=\omega^{\prime}_{n}$ for all $n<N$). Denote the (continuous) left shift by $\sigma:\Omega\to\Omega$. Equip the product $S:=\Omega\times\mathbb{P}(\mathcal{H})$ with the product metric
$d_{S}\bigl((\omega,[u]),(\omega^{\prime},[u^{\prime}])\bigr)=d_{\Omega}(\omega,\omega^{\prime})+d_{FS}([u],[u^{\prime}]).$
Set
$\mathcal{D}=\{(\omega,[u])\in S:[u]\in D_{\omega_{0}}\}.$
Definition 3.2.
Define the skew–product map $F:\mathcal{D}\to S$ by
$F(\omega,[u])=(\sigma\omega,f_{\omega_{0}}([u])).$
Here $\omega_{0}$ denotes the zeroth coordinate of $\omega$ (the symbol corresponding to time 0).
Note that, in general, $F$ is not a map from all of $\Omega\times\mathbb{P}(\mathcal{H})$ to itself; and it is continuous because each fibre map $f_{j}$ is continuous on $D_{j}$ and $\sigma$ is continuous.
Definition 3.3.
A selector is any function
$D:\mathbb{P}(\mathcal{H})\to\Omega,\qquad[u]\mapsto D([u])=(D([u])_{n})_{n\in\mathbb{N}_{0}}.$
No regularity is assumed a priori: $D$ may be completely arbitrary. We will come back to the physical motivation for considering such general maps. We denote by $\mathfrak{D}$ the set of all possible selectors.
Definition 3.4.
Let
$T:U(\mathcal{H})\times\mathfrak{D}\to\mathbb{P}(\mathcal{H})^{\mathbb{P}(\mathcal{H})}$
be the map defined by
$(U,D)\mapsto T(U,D):\mathbb{P}(\mathcal{H})\to\mathbb{P}(\mathcal{H}),$
where the induced map is defined as
$T(U,D)([u]):=f_{D([u])_{0}}([u]),$
whenever the right-hand side is well defined.
We may denote the induced map $T(U,D)$ by $T_{U,D}$, or simply $T$ when the unitary operator $U$ and the selector $D$ are clear from the context.
Under such a general definition, it is clear that not every selector makes physical sense. In the following, we will see some necessary criteria that a selector map has to meet to be considered physically meaningful.
First of all, we have to ensure that the stories of collapse events seen by distinct states are consistent with each other. This is accomplished by the following requirement.
Definition 3.5.
A selector $D:\mathbb{P}(\mathcal{H})\to\Omega$ is said to be compatible with the skew–product dynamics if, for every $u]\in\mathbb{P}(\mathcal{H})$ for which $T([u])$ is defined, the following identity holds:
$D\bigl(T([u])\bigr)=\sigma\bigl(D([u])\bigr).$ (2)
Moreover, we have to ensure that collapse events at each point produce observable outcomes only having positive probability.
This motivates the following request.
Definition 3.6.
A selector $D:\mathbb{P}(\mathcal{H})\to\Omega$ is called admissible if $D([u])_{0}\neq j\text{ whenever }[u]\notin D_{j}\text{ (or equivalently $Uu\in\ker(P_{j})$)}$.
Then the map $T([u])=f_{D([u])_{0}}([u])$ is well defined for every $u]\in\mathbb{P}(\mathcal{H})$.
Definition 3.7.
A realization map is a selector that is compatible and admissible.
It is easily seen that such a map exists.
Let us observe that, in a general open quantum system, none of the measurement channels (including the “blank” no-collapse case) can be regarded as strictly forbidden in any open region of the state space. More precisely, every outcome label $j\in\Lambda$ must remain topologically accessible in arbitrarily small neighborhoods of every pure state. Therefore, each outcome label $j\in\Lambda$ occurs arbitrarily close to every pure state, which leads to the topological condition
$\overline{\{[u]\in\mathbb{P}(\mathcal{H}):D([u])_{0}=j\}}=\mathbb{P}(\mathcal{H})\qquad\text{for all }j\in\Lambda.$ (3)
Note that we are not assuming (3), but we stress that we are allowing it. Stronger formulations (for instance, requiring a positive Fubini–Study measure or open support of each outcome set) would impose quantitative rather than merely topological versions of the same idea, but note that, even under this minimal topological condition, the map $T_{U,D}$ is discontinuous everywhere.
Lemma 3.8.
If $P_{j}$ is not the zero operator for $j\neq 0$, the sets
$D_{j}=\{[u]\in\mathbb{P}(\mathcal{H}):\ \|P_{j}Uu\|\neq 0\}$
are open for all $j\in\Lambda$, and $D_{j}$ is dense for every $j\neq 0$. In particular $D_{0}=\mathbb{P}(\mathcal{H})$.
Proof.
Openness follows from continuity of $u]\mapsto\|P_{j}Uu\|$. If $j\neq 0$, then $P_{j}U\neq 0$, so $\ker(P_{j}U)$ is a proper subspace of $\mathcal{H}$.
Consequently, the projectivization $\mathbb{P}(\ker(P_{j}U))$ is a proper (closed) projective subspace of $\mathbb{P}(\mathcal{H})$, hence it has empty interior. Therefore,
$D_{j}=\mathbb{P}(\mathcal{H}\setminus\mathbb{P}(\ker(P_{j}U))$
is dense. Finally, $P_{0}=I$ gives $D_{0}=\mathbb{P}(\mathcal{H})$. ∎
From now on we fix a realization map $D:\mathbb{P}(\mathcal{H})\to\Omega$ and read it as an outcome itinerary: at step $n$, i.e. if a measurement of the fixed PVM $\{P_{j}\}_{j\in\Lambda}$ is performed, the symbol we use is $D([u])_{n}$. We analyze the dynamics conditional on this single branch.
Remark 3.9.
Note that the existence of a realization map does not conflict with the Kochen–Specker (KS) no-go theorem ([kochen2011problem]). In fact, it addresses a completely different point. KS forbids noncontextual, dispersion-free value assignments to all projections simultaneously (in dimension $\geq 3$), preserving functional relations across incompatible observables. By contrast, throughout we fix a single measurement context $\{P_{j}\}_{j\in\Lambda}$ (one PVM), and the realization map $D:\mathbb{P}(\mathcal{H})\to\Omega$ assigns an outcome sequence only for this context. The assignment is explicitly contextual/history-dependent and makes no claims about outcomes for other, incompatible PVMs. Hence $D$ is not a global valuation on the projection lattice, and the hypotheses of KS are not met.
Summarizing, the map $T_{U,D}$ behaves as follows: if $D([u])=\omega_{0}\omega_{1}\omega_{2}\dots\in\Omega$ then
$T([u])=f_{\omega_{0}}([u]),$
provided that $u]\in D_{\omega_{0}}$ so that the expression makes sense. Thus, $T$ is obtained by applying to $[Uu]$ the fiber map corresponding to the first symbol in its realization sequence. Notice that $T$ is defined on the whole space $\mathbb{P}(\mathcal{H})$ if and only if the selector $D$ is admissible.
Remark 3.10.
We make no non-degeneracy assumption on eigenspaces. However, even if all nonblank projections $P_{j}$ ($j\neq 0$) have rank 1 (so each $f_{j}$ collapses to a single point of $\mathbb{P}(\mathcal{H})$), the presence of the “blank” outcome 0 with $P_{0}=I$ interleaves unitary motion with discontinuous point–collapses. Under the density property 3, $D(\cdot)_{0}$ is nowhere locally constant, which forces $T$ to be nowhere continuous. Hence our chain–recurrence conclusion cannot be deduced trivially even in the rank–one collapse case.
4 Reversibility
Definition 4.1 (Integrated energetic cost).
Let $H$ be the ambient Hamiltonian and let a bounded self-adjoint perturbation $\Delta H(t)$ act on the purely unitary segments between collapse events, so that the evolution obeys to:
$i\,\partial_{t}W(t)=(H+\Delta H(t))W(t),$
for a suitable unitary operator $W$. For a time interval $I\subset\mathbb{R}$, the integrated energetic cost of $\Delta H$ on $I$ is
$\mathcal{E}_{I}[\Delta H]\;:=\;\int_{I}\!\|\Delta H(t)\|_{*}\,dt,$
where $\|\cdot\|_{*}$ denotes any unitarily invariant matrix norm (by default, the operator norm).
Definition 4.2.
Let $\gamma$ be a finite forward itinerary given by:
$\gamma=\big([u_{0}],[u_{1}],\dots,[u_{N}]\big),\qquad[u_{k+1}]=T_{W_{k}(k+1,k),D}([u_{k}]),$
where for every $k=0,\ldots,N-1$ let $\Delta H_{k}(t)$ be a bounded self-adjoint perturbation acting on the unitary segments $I_{k}:=[k,k+1]$ between collapse events, so that the evolution obeys $i\,\partial_{t}W_{k}(t)=(H+\Delta H_{k}(t))W_{k}(t)$ for a suitable unitary operator $W_{k}$. Then, the energetic cost of the itinerary $\gamma$ is defined as:
$\mathcal{E}_{D}(\gamma)=\sum_{k=0}^{N-1}\mathcal{E}_{I_{k}}[\Delta H_{k}],$
where $\mathcal{E}_{I_{k}}[\Delta H_{k}]$ is the integrated energetic cost of $\Delta H_{k}$ on $I_{k}$.
For $\epsilon>0$, we denote by $B_{\epsilon}([Uu])$ the open $d_{FS}$-ball of radius $\epsilon$ centered at $[Uu]$ in $\mathbb{P}(\mathcal{H})$.
Let us prove the following Lemma:
Lemma 4.3.
Let $u\in\mathcal{H}$ and $0<\tau_{0}\leq\tau<\tau_{1}$. For every $\epsilon>0$ and $v]\in B_{\epsilon}([U(\tau_{1},\tau_{0})u])$, there exist a rank-$2$ skew-Hermitian operator $K(v,\delta)$ with $\|K(v,\delta)\|=O(\epsilon)$ and a family of self-adjoint perturbations
$\Delta H_{\epsilon}(t)\;=\;U(t,\tau)\,\widetilde{H}\,U(t,\tau)^{\dagger}\qquad t\in[\tau,\tau_{1}],$
[…]
$V(t,\tau)=\mathcal{T}\exp\!\Big(-i\int_{\tau}^{t}\big(H+\Delta H_{\epsilon}(s)\big)\,ds\Big)$
satisfies
$V(\tau_{1},\tau)\,U(\tau,\tau_{0})u=v,\qquad\text{and}\qquad\lim_{\epsilon\to 0}\|V(\tau_{1},\tau)-U(\tau_{1},\tau)\|=0.$
(Notice that no commutation relation between $H$ and $K(v,\delta)$ is assumed.)
Proof.
Set $w:=U(\tau_{1},\tau_{0})u$. Without loss of generality, let us assume $\langle w,v\rangle\geq 0$ (replacing $v\mapsto e^{-i\arg\langle w,v\rangle}v$ if necessary) in such a way to preserve $\|v\|=1$ and $d_{\mathrm{FS}}([w],[v])$. Let
$\delta:=d_{\mathrm{FS}}([w],[v])=\arccos\langle w,v\rangle\leq\epsilon$
and define $K(v,\delta)$ on $\operatorname{span}\{w,v\}$ by
$K(v,\delta):=\frac{\arccos|\langle v,w\rangle|}{\sqrt{1-|\langle v,w\rangle|^{2}}}\Big(|v\rangle\langle w|-|w\rangle\langle v|\Big).$
The rank-2 operator $K(v,\delta)$ is clearly skew-Hermitian and it is a rotational generator of angle $\delta$ on the plane, i.e. it verifies $e^{K(v,\delta)}w=v.$ Moreover, $\|K(v,\delta)\|=\delta=O(\epsilon)$. Let us now set the following
$R$ $\coloneqq e^{K(v,\delta)}$
$\Delta\tau$ $\coloneqq\tau_{1}-\tau>0$
[…]
$\Delta H_{\epsilon}(t)$ $\coloneqq U(t,\tau)\,\widetilde{H}\,U(t,\tau)^{\dagger},\qquad t\in[\tau,\tau_{1}].$
Let $V(t,\tau)$ solve the Schrödinger evolution equation
$\partial_{t}V(t,\tau)=-i(H+\Delta H_{\epsilon}(t))V(t,\tau)$
with $V(\tau,\tau)=I$, and define
$X(t):=U(t,\tau)^{\dagger}\,V(t,\tau).$
Then
$\,\partial_{t}X(t)$ $=iU(t,\tau)^{\dagger}\,H\,V(t,\tau)-iU(t,\tau)^{\dagger}\,(H+\Delta H_{\epsilon}(t))\,V(t,\tau)=$
$=-iU(t,\tau)^{\dagger}\,\Delta H_{\epsilon}(t)\,V(t,\tau)=-iU(t,\tau)^{\dagger}\,\Delta H_{\epsilon}(t)\,U(t,\tau)\,X(t)=-i\widetilde{H}\,X(t).$
where the last equality follows by the fact that $U(t,\tau)^{\dagger}\,\Delta H_{\epsilon}(t)\,U(t,\tau)=\widetilde{H}$. The condition $X(\tau)=I$ implies that $X(t)=e^{-i\widetilde{H}(t-\tau)}$ and then
$X(\tau_{1})$ $=\exp\!\big(\!-i\widetilde{H}\Delta\tau\big)=\exp\!\Big(\!-i\frac{i}{\Delta\tau}\log\Big(e^{iH\Delta\tau}\,R\,e^{-iH\Delta\tau}\Big.)\Delta\tau\Big)=$
$=e^{iH\Delta\tau}\,R\,e^{-iH\Delta\tau}=U(\tau_{1},\tau)^{\dagger}\,R\,U(\tau_{1},\tau).$
Therefore
$V(\tau_{1},\tau)=U(\tau_{1},\tau)\,X(\tau_{1})=U(\tau_{1},\tau)\,U(\tau_{1},\tau)^{\dagger}\,R\,U(\tau_{1},\tau)=R\,U(\tau_{1},\tau),$
and
$V(\tau_{1},\tau)\,U(\tau,\tau_{0})\,u=R\,U(\tau_{1},\tau)\,U(\tau,\tau_{0})\,u=R\,w=v.$
It remains to prove that
$\lim_{\epsilon\to 0}\|V(\tau_{1},\tau)-U(\tau_{1},\tau)\|=0.$
First observe that, since $\|K(v,\delta)\|\to 0$ as $\epsilon\to 0$, we have $e^{K(v,\delta)}\to I$ in operator norm. Now choose a rank-2 Hermitian $J_{S}$ supported on $\operatorname{span}\{w,v\}$ with $\|J_{S}\|=1$ such that we can write $R=e^{K(v,\delta)}=e^{-i\delta J_{S}}$. Then we have that
$\widetilde{H}$ $=\frac{i}{\Delta\tau}\log\!\big(U(\tau_{1},\tau)^{\dagger}R\,U(\tau_{1},\tau)\big)=\frac{\delta}{\Delta\tau}\,J_{c}\qquad\text{where}\qquad J_{c}:=U(\tau_{1},\tau)^{\dagger}J_{S}\,U(\tau_{1},\tau).$
Therefore,
$\sup_{t\in[\tau,\tau_{1}]}\!\|\Delta H_{\epsilon}(t)\|=\|\widetilde{H}\|=\frac{\delta}{\Delta\tau}\qquad\text{and}\qquad\mathcal{E}_{\Delta\tau}[\Delta H_{\epsilon}]=\int_{\tau}^{\tau_{1}}\!\|\Delta H_{\epsilon}(t)\|\,dt=\Delta\tau\cdot\frac{\delta}{\Delta\tau}=\delta,$ (4)
which means that the instantaneous bound vanishes as $\delta\to 0$ for fixed $\Delta\tau$ and the integrated energetic cost is exactly $\delta$.
Moreover, from $\partial_{t}U(t,\tau)=-iHU(t,\tau)$ and $\partial_{t}V(t,\tau)=-i(H+\Delta H_{\epsilon}(t))V(t,\tau)$ we have that $V(t,\tau)-U(t,\tau)$ satisfies the following differential equation:
$\partial_{t}\big(V(t,\tau)-U(t,\tau)\big)=-iH\big(V(t,\tau)-U(t,\tau)\big)-i\Delta H_{\epsilon}(t)V(t,\tau),$
with solution
$V(t,\tau)-U(t,\tau)=-i\int_{\tau}^{t}U(t,s)\Delta H_{\epsilon}(s)V(s,\tau)\ ds.$
Finally, exploiting the result in norm, we have the following inequality
$\|V(t,\tau)-U(t,\tau)\|\leq\int_{\tau}^{t}\|\Delta H_{\epsilon}(s)\|\ ds.$
By (4), it follows that $\|V(\tau_{1},\tau)-U(\tau_{1},\tau)\|\leq\delta\leq\epsilon,$ from which the claim follows. ∎
Remark 4.4.
[…]
The following result is key to our reasoning. For the proof, see Appendix. The proof is a refinement of that of Theorem 3.9 in [della2025chains]. To make the present work independent, we cover it in detail without relying on the proof of the weaker result.
Theorem 4.5.
Assume that $D$ is an admissible and compatible selector. Then the dynamical system $(\mathbb{P}(\mathcal{H}),T_{U,D})$ has a strongly chain-recurrent point and a topologically closed, invariant, strongly chain-transitive subsystem.
The following result establishes that we can construct chains in $(\mathbb{P}(\mathcal{H}),T_{U,D})$ with unitary perturbations having arbitrary small energetic cost.
Theorem 4.6.
Consider the dynamical system $(\mathbb{P}(\mathcal{H}),T_{U,D})$. For every pair $u],[v]\in\mathbb{P}(\mathcal{H})$ with $u]\,\mathcal{SC}_{d}\,[v]$ and every $\epsilon>0$, there exists a finite family of self-adjoint, time-dependent perturbations $\{\Delta H_{\epsilon_{k}}(t)\}_{k=1}^{N-1}$ satisfying
$\sum_{k=1}^{N-1}\mathcal{E}_{[\tau_{k}^{-},\tau_{k}^{+}]}[\Delta H_{\epsilon_{k}}]=\;O\!\Big(\sum_{k=1}^{N-1}\epsilon_{k}\Big)\;=\;O(\epsilon)\qquad(\tau_{k}^{-}:=k+\tfrac{1}{3},\ \tau_{k}^{+}:=k+\tfrac{2}{3})$
and such that $d_{FS}(\widetilde{T}\,[u],[v])<\epsilon$, where
$\widetilde{T}\,[u]:=T_{W(N-1,N-2),D}\big(T_{W(N-2,N-3),D}\big(\ldots\big(T_{W(1,0),D}[u]\big)\ldots\big).$
The total evolution over $N$ steps, with perturbations applied at intervals $\tau^{-}_{k},\tau^{+}_{k}$] $(k=1,\ldots,N-1)$, is given by:
$W(t)\;=\;\mathcal{T}\exp\!\Big(-i\!\int_{0}^{t}H_{\mathrm{tot}}(s)\,ds\Big)\qquad\text{with}\qquad H_{\mathrm{tot}}(t):=H+\sum_{k=1}^{N-1}\mathbf{1}_{[\tau_{k}^{-},\tau_{k}^{+}]}(t)\,\Delta H_{\epsilon_{k}}(t).$
Proof.
Fix $\epsilon>0$. Since $u]\,\mathcal{SC}_{d}\,[v]$, there is a strong $\epsilon$-chain
$u_{0}]=[u],[u_{1}],\dots,[u_{N-1}],[u_{N}]=[v]$
with $d_{FS}(T_{U,D}[u_{k-1}],[u_{k}])<\epsilon_{k}$ for $k=1,\dots,N$ and $\sum_{k=1}^{N}\epsilon_{k}<\epsilon$.
For every $k=1,\dots,N-1$ we set
$w_{k}]:=T_{U,D}[u_{k-1}],\quad\tau_{k}^{-}:=k+\tfrac{1}{3},\quad\tau_{k}^{+}:=k+\tfrac{2}{3},\quad\Delta\tau_{k}:=\tau_{k}^{+}-\tau_{k}^{-}$
For each $k=1,\dots,N-1$, apply Lemma 4.3 with
$\tau_{0}=k,\quad\tau=\tau_{k}^{-},\quad\tau_{1}=\tau_{k}^{+},\quad u=w_{k},\quad v=U(\tau_{k}^{+},k)\,u_{k},$
to obtain a perturbation supported on the interval $\tau_{k}^{-},\tau_{k}^{+}]$ of the form
$\Delta H_{\epsilon_{k}}(t)\;=\;U(t,\tau_{k}^{-})\,\widetilde{H}_{k}\,U(t,\tau_{k}^{-})^{\dagger}\qquad t\in[\tau_{k}^{-},\tau_{k}^{+}],$
[…]
$V_{k}(\tau_{k}^{+},\tau_{k}^{-}):=\mathcal{T}\exp\!\Big(-i\!\int_{\tau_{k}^{-}}^{\tau_{k}^{+}}\big(H+\Delta H_{\epsilon_{k}}(s)\big)\,ds\Big)$
then satisfies
$V_{k}(\tau_{k}^{+},\tau_{k}^{-})\,U(\tau_{k}^{-},k)\,w_{k}\;=\;U(\tau_{k}^{+},k)\,u_{k},\qquad\|V_{k}(\tau_{k}^{+},\tau_{k}^{-})-U(\tau_{k}^{+},\tau_{k}^{-})\|\xrightarrow[\epsilon_{k}\to 0]{}0.$
In particular, if we set $W_{k}:=U(k+1,\tau_{k}^{+})\,V_{k}(\tau_{k}^{+},\tau_{k}^{-})\,U(\tau_{k}^{-},k)$, we find
$W_{k}\,w_{k}=U(k+1,\tau_{k}^{+})U(\tau_{k}^{+},k)\,u_{k}=U(k+1,k)\,u_{k}.$
Therefore,
$T_{W_{k}(k+1,k),D}[w_{k}]=T_{U,D}[u_{k}]=[w_{k+1}].$ (5)
Define the total evolution over the time-interval $[0,N]$ by:
$W(t)=\mathcal{T}\exp\!\big(-i\!\int_{0}^{t}H_{\mathrm{tot}}(s)\,ds\big)$
with $H_{\mathrm{tot}}(t)$ as defined in the statement. If we take $t=N$, we can equivalently write the total evolution with perturbations applied at intervals $\tau^{-}_{k},\tau^{+}_{k}$] $(k=1,\ldots,N-1)$ as the left-ordered product:
$W(N)\;=\;\overleftarrow{\prod_{k=1}^{N-1}}\Big[W_{k}(k+1,k)\Big]\;U(1,0)=\;\overleftarrow{\prod_{k=1}^{N-1}}\Big[\,U(k+1,\tau_{k}^{+})\,V_{k}(\tau_{k}^{+},\tau_{k}^{-})\,U(\tau_{k}^{-},k)\,\Big]\;U(1,0).$
Since $W(k+1,k)=W_{k}(k+1,k)$ for $k=1,\ldots,N-1$, by (5) it follows that
$T_{W(k,k-1),D}\big(T_{W(k-1,k-2),D}\big(\ldots\big(T_{W(1,0),D}[u]\big)\ldots\big)=T_{U,D}[u_{k-1}]$
for every $k=1,\ldots,N-1$. Therefore,
$d_{FS}(\widetilde{T}\,[u],[v])=d_{FS}(T_{U,D}[u_{N-1}],[v])<\epsilon_{N}<\epsilon.$ (6)
Notice that $W(N)u=Uu_{N-1}$ if $D([u_{k}])_{0}=0$ for every $k=0,\ldots,N-1$.
Finally, the energetic cost follows from the unitary invariance of the operator norm:
$\mathcal{E}_{\Delta\tau_{k}}[\Delta H_{\epsilon_{k}}]=\int_{\tau_{k}^{-}}^{\tau_{k}^{+}}\!\|\Delta H_{\epsilon_{k}}(t)\|\,dt=O(\epsilon_{k}),$
hence $\sum_{k=1}^{N-1}\mathcal{E}_{\Delta\tau_{k}}[\Delta H_{\epsilon_{k}}]=\sum_{k=1}^{N-1}\int\|\Delta H_{\epsilon_{k}}\|=O(\sum_{k}\epsilon_{k})=O(\epsilon)$. ∎
5 Conclusion
Theorems 4.5 and 4.6 tell us the following fact: in every finite-dimensional, open quantum system, there is a topologically closed, invariant subsystem $S$ where any two states are connected via pseudo-orbits with arbitrarily small Fubini-Study error and energetic cost. By Def. 1.1, this means that $S$ is operationally reversible: no operational arrow of time arises.
It is perhaps useful to frame this result by analogy with the logical structure of classical thermodynamics. The second law provides a universal lower bound on irreversibility, while Carnot’s theorem is an existence result showing that this bound can in fact be saturated: it identifies ideal cycles that achieve reversibility under suitable structural constraints. Landauer’s principle plays a similar role to the second law in the informational setting, furnishing a universal lower bound on dissipation for logically irreversible operations. Our result stands, structurally, in the same relation to Landauer’s principle as Carnot’s theorem does to the second law: it is an existence statement identifying conditions (finite dimensionality and compactness) under which the bound is asymptotically saturated. In this sense, operational quasi-reversibility is enforced within invariant subsystems, in the absence of information erasure or non-compact limits.
Let us finally remark that our argument crucially relies on the retention of the infinite record of outcomes provided by the realization map. This infinite history is essential because, to steer the system toward a target state with arbitrarily small FS error $\epsilon>0$, the perturbations must be tailored to the specific sequence of realized collapse outcomes encountered along the chain. Although only finitely many corrections are required to achieve any fixed $\epsilon$-precision, the number and nature of these corrections generally become unbounded as $\epsilon\to 0$, necessitating, in principle, access to the entire future itinerary to prescribe the appropriate steering strategy at every scale. This echoes Bennett’s resolution of Maxwell’s demon paradox, where he showed that a demon’s finite memory forces logically irreversible erasure to accommodate new data, thereby incurring an entropic cost that preserves the second law ([bennett1982thermodynamics]). Our no-erasure framework mirrors the idealized demon scenario, where an infinite memory sidesteps dissipation. Yet, as Bennett argued for any physical demon, real-world implementations face finite-size constraints that inevitably lead to erasure. Analogously, any practical embodiment of our dynamics with a truncated record would introduce operational irreversibility, which is of course consistent with the second law.
Appendix
In this Appendix we address the proof of Theorem 4.5.
We start recalling the following result that shows that in a compact dynamical system, without any regularity assumption on the map, there always exists a generalized recurrent point for the system, that is, a point $x\in X$ such that $x\in\bigcap_{d\in M}\mathcal{SCR}_{d}(f)=:GR(f)$ where $M$ is the set of the metrics compatible with the topology of $X$.
Lemma 5.1.
[della2025chains, Theorem 3.5] Let $(X,f)$ be a compact dynamical system. Then $GR(f)\neq\emptyset$.
The starting step in the proof of the previous result is the construction, by transfinite induction, of a sequence of points $\{x_{\alpha}\}$ described below.
Assume that every point in the space $X$ has an infinite orbit, otherwise the claim of Theorem 5.1 follows easily. Pick $x_{0}:=x\in X$ and set $x_{k}:=f(x_{k-1})$ for all $k\in\mathbb{N}$. Take $y\in\mathcal{O}^{\prime}(x)$, which is nonempty due to the compactness of $X$, and set $x_{\omega}:=y$. Let $\alpha>\omega$ be an ordinal number and assume that $x_{\beta}$ has been defined for every $\beta<\alpha$. Define the following sets:
$S_{\beta,\alpha}:=\{x_{\eta}\}_{\beta\leq\eta<\alpha}\quad,\quad S_{\alpha}(x):=S_{0,\alpha}(x).$ (7)
If $\alpha$ is a successor ordinal, we set $x_{\alpha}:=f(x_{\alpha-1})$. Otherwise, if $\alpha$ is a limit ordinal, we set
$S^{*}_{\alpha}:=\bigcap_{\beta<\alpha}(S_{\beta,\alpha}(x))^{\prime},$ (8)
and $x_{\alpha}:=y$ with $y$ any point in $S^{*}_{\alpha}(x)$. Notice that $S^{*}_{\alpha}(x)\neq\emptyset$ (see, e.g. [munkres2013topology, Th. 26.9, p. 169]), since $|S_{\beta,\alpha}(x)|\geq\aleph_{0}$ for every $\beta<\alpha$ so that $(S_{\beta,\alpha}(x))^{\prime}$ is a closed and nonempty set. The second step in the proof of [della2025chains, Theorem 3.5] is to show that for every ordinal $\alpha$,
$x_{\beta}\,\mathcal{SC}_{d}\,x_{\eta}\quad\text{whenever}\quad 0\leq\beta<\eta\leq\alpha.$ (9)
Finally, in the last step of the proof, it is shown, by a cardinality argument, that there exists an ordinal $\alpha$ such that $x_{\alpha}=x_{\beta}$ for some $\beta<\alpha$, and thus $x_{\alpha}\in GR(f)$.
Moreover, by exploiting the proof of property (9) in [della2025chains] we have the following
Remark 5.2.
For every $\beta<\eta\leq\alpha$ and every $\epsilon>0$, there is a strong $(\epsilon,d)$-chain from $x_{\beta}$ to $x_{\eta}$ whose points belong to $S_{\beta,\eta+1}(x)$. Therefore, if $x_{\alpha}=x_{\beta}$ for some $\alpha>\beta$, then by the transitivity of $\mathcal{SC}_{d}$ we have that the set $S_{\beta,\alpha}(x)$ is internally $\mathcal{SC}_{d}$-transitive, that is, for every $y,z\in S_{\beta,\alpha}(x)$ and for every $\epsilon>0$ there exists a strong $(\epsilon,d)$-chain from $y$ to $z$ whose points belong to $S_{\beta,\alpha}(x)$.
The following lemma is a direct consequence of [della2025chains, Lemma 3.6 and Remark 3.8] and Remark 5.2.
Lemma 5.3.
Let $(X,f)$ be a compact dynamical system. Then $X$ has an invariant, internally $\mathcal{SC}_{d}$-transitive subset. ∎
In particular, by Lemma 5.1, we can say that for $x\in X$, there exists a least ordinal $\alpha>0$ such that $x_{\alpha}\in S_{\alpha}(x)$ and so, in particular, $x_{\alpha}=x_{\beta}$ for some $\beta<\alpha$. Moreover, the set $S_{\beta,\alpha}(x)$ is an invariant and internally $\mathcal{SC}_{d}$-transitive subset of $X$.
In [della2025chains] it is shown that a compact dynamical system has a chain-transitive subsystem. Using the previous lemma and Remark 5.2, we want to prove that every compact dynamical system has, in particular, a $\mathcal{SC}_{d}$-transitive subsystem.
Now are ready to prove the result, that we formulate here in the abstract topological dynamical notation.
Theorem 5.4.
Let $(X,f)$ be a compact dynamical system. Then $(X,f)$ has a $\mathcal{SC}_{d}$-transitive subsystem.
Proof.
We need to show that there is a closed, invariant and internally $\mathcal{SC}_{d}$-transitive subset of $X$. Suppose, towards a contradiction, that $(X,f)$ does not have a $\mathcal{SC}_{d}$-transitive subsystem. In particular, we can assume that every orbit is infinite. Let $x^{0}\in X$. By Lemma 5.1, there exists a least ordinal $\alpha_{0}>0$ such that $x^{0}_{\alpha_{0}}\in S_{\alpha_{0}}(x^{0})$, and so $x^{0}_{\alpha_{0}}=x^{0}_{\beta_{0}}$ for some $\beta_{0}<\alpha_{0}$. By Lemma 5.3 and Remark 5.2, the set $S_{\beta_{0},\alpha_{0}}(x^{0})$ is an invariant, internally $\mathcal{SC}_{d}$-transitive subset. Since $S_{\beta_{0},\alpha_{0}}(x^{0})$ cannot be closed, there exists some $y\in(S_{\beta_{0},\alpha_{0}}(x^{0}))^{\prime}\setminus S_{\beta_{0},\alpha_{0}}(x^{0})$. Let $x\in S_{\alpha_{0}}(x^{0})$ and $\epsilon>0$, we want to show that there exists a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{y\}$. Let $\beta_{0}\leq\eta<\alpha_{0}$ be such that $d(x_{\eta},y)<\epsilon/2$. By property (9) and Remark 5.2, it follows that there exists a strong $(\epsilon/2,d)$-chain $x=y_{0},y_{1},\ldots,y_{n-1},y_{n}=x_{\eta}$ from $x$ to $x_{\eta}$ whose points belong to $S_{\alpha_{0}}(x^{0})$. By the triangle inequality,
$\sum_{i=0}^{n-2}d(f(y_{i}),y_{i+1})+d(f(y_{n-1}),y)\leq\sum_{i=0}^{n-1}d(f(y_{i}),y_{i+1})+d(x_{\eta},y)<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon$ (10)
then, the points $x=y_{0},y_{1},\ldots,y_{n-1},y$ form a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{y\}$.
If $y\neq x^{0}$, we set $x^{1}:=y$. Otherwise, if $y=x^{0}$, it follows that $S_{\alpha_{0}}(x^{0})$ is an invariant and an internally $\mathcal{SC}_{d}$-transitive subset. Since it cannot be a closed set, there is some $z\in(S_{\alpha_{0}}(x^{0}))^{\prime}\setminus S_{\alpha_{0}}(x^{0})$, and we set $x^{1}:=z\,(\neq x^{0})$. With an analogous argument used in (10), we find $S_{\alpha_{0}}(x^{0})\,\mathcal{SC}_{d}\,x^{1}$, and for every $x\in S_{\alpha_{0}}(x^{0})$ and for every $\epsilon>0$ there exists a strong $(\epsilon,d)$-chain from $x$ to $x^{1}$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{x^{1}\}$.
In summary, we have shown that the set $\{x^{0},x^{1}\}$ has the following properties:
  • (a1)
    there exists a least ordinal $\alpha_{0}$ such that $x^{0}_{\alpha_{0}}=x^{0}_{\beta_{0}}$ for some $\beta_{0}<\alpha_{0}$ and $S_{\beta_{0},\alpha_{0}}(x^{0})$ is an invariant, internally $\mathcal{SC}_{d}$-transitive subset;
  • (b1)
    $S_{\alpha_{0}}(x^{0})\,\mathcal{SC}_{d}\,x^{1}$, and for every $x\in S_{\alpha_{0}}(x^{0})$ and for every $\epsilon>0$ there exists a strong $(\epsilon,d)$-chain from $x$ to $x^{1}$ whose points belong to $S_{\alpha_{0}}(x^{0})\cup\{x^{1}\}$;
  • (c1)
    $x^{0}\neq x^{1}$.
Let us proceed now by transfinite induction. Let $\lambda>1$ be an ordinal number. Assume that $x^{\gamma}$ has been defined for every $0\leq\gamma<\lambda$ and that the set $\{x^{\gamma}\}_{\gamma<\lambda}$ has the following property:
  1. (aλ)
    for every $0\leq\gamma<\lambda$, there exists a least ordinal $\alpha_{\gamma}$ such that $x^{\gamma}_{\alpha_{\gamma}}=x^{\gamma}_{\beta_{\gamma}}$ for some $\beta_{\gamma}<\alpha_{\gamma}$ and $S_{\beta_{\gamma},\alpha_{\gamma}}(x^{\gamma})$ is an invariant, internally $\mathcal{SC}_{d}$-transitive subset.
Moreover, setting for every $\beta<\alpha\leq\lambda$
$M_{\beta,\alpha}:=\bigcup_{\beta\leq\gamma<\alpha}S_{\alpha_{\gamma}}(x^{\gamma}),$
assume also the following properties:
  1. (bλ)
    For every $\eta<\xi<\lambda$, we have $S_{\alpha_{\eta}}(x^{\eta})\,\mathcal{SC}_{d}\,x^{\xi}$, and for every $x\in S_{\alpha_{\eta}}(x^{\eta})$ and every $\epsilon>0$ there exists a strong $(\epsilon,d)$-chain from $x$ to $x^{\xi}$ whose points belong to $M_{\eta,\xi}\cup\{x^{\xi}\}$;
  2. (cλ)
    $x^{\eta}\neq x^{\xi}$ whenever $\eta\neq\xi$ with $\eta,\xi<\lambda$ .
We say that a point $z$ verifies property Pλ if, for every $\eta<\lambda$, we have $S_{\alpha_{\eta}}(x^{\eta})\,\mathcal{SC}_{d}\,z$, and for every $x\in S_{\alpha_{\eta}}(x^{\eta})$ and every $\epsilon>0$, there exists a strong $(\epsilon,d)$-chain from $x$ to $z$ whose points belong to $M_{\eta,\lambda}\cup\{z\}$. Notice that, assuming (bλ), to prove property (bλ+1) it is sufficient to show that the point $x^{\lambda}$ verifies property Pλ. We now proceed with the definition of the point $x^{\lambda}$.
  • Case 1.
    $\lambda$ is a successor ordinal.
    Consider the set $S_{\beta_{\lambda-1},\alpha_{\lambda-1}}(x^{\lambda-1})$, which is invariant and internally $\mathcal{SC}_{d}$-transitive by property (aλ). Since it cannot be a closed set, there exists some
    $y\in(S_{\beta_{\lambda-1},\alpha_{\lambda-1}}(x^{\lambda-1}))^{\prime}\setminus S_{\beta_{\lambda-1},\alpha_{\lambda-1}}(x^{\lambda-1}).$
    The point $y$ verifies property Pλ. Indeed, with an analogous argument used in (10), we have that for every $x\in S_{\alpha_{\lambda-1}}(x^{\lambda-1})$ and for every $\epsilon>0$ there exists a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $S_{\alpha_{\lambda-1}}(x^{\lambda-1})\cup\{y\}$. Fix $\eta<\lambda$ and $\epsilon>0$ and take $x\in S_{\alpha_{\eta}}(x^{\eta})$. Let $x^{\lambda-1}=y_{0},y_{1},\ldots,y_{n-1},y_{n}=y$ ($n\in\mathbb{N}$) be a strong $(\epsilon/2,d)$-chain from $x^{\lambda-1}\in S_{\alpha_{\lambda-1}}(x^{\lambda-1})$ to $y$ whose points belong to $S_{\alpha_{\lambda-1}}(x^{\lambda-1})\cup\{y\}$. By property (bλ), let $x=z_{0},z_{1},\ldots,z_{m-1},z_{m}=x^{\lambda-1}$ ($m\in\mathbb{N}$) be a strong $(\epsilon/2,d)$-chain from $x$ to $x^{\lambda-1}$ whose points belong to $M_{\eta,\lambda-1}\cup\{x^{\lambda-1}\}$. Therefore, noting that $M_{\eta,\lambda}=M_{\eta,\lambda-1}\cup S_{\alpha_{\lambda-1}}(x^{\lambda-1})$, we have that $x=z_{0},z_{1},\ldots,z_{m-1},x^{\lambda-1},y_{1},\ldots,y_{n-1},y_{n}=y$ is a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $M_{\eta,\lambda}\cup\{y\}$. Indeed, since $z_{m}=x^{\lambda-1}=y_{0}$, we have
    $\sum_{i=0}^{m-2}d(f(z_{i}),z_{i+1})+d(f(z_{m-1}),x^{\lambda-1})+d(f(x^{\lambda-1}),y_{1})+\sum_{i=0}^{n-1}d(f(y_{i}),y_{i+1})=$ (11)
    $=\sum_{i=0}^{m-1}d(f(z_{i}),z_{i+1})+\sum_{i=0}^{n-1}d(f(y_{i}),y_{i+1})<\frac{\epsilon}{2}+\frac{\epsilon}{2}=\epsilon.$
    Consider the following two cases.
    • If $y\notin\{x^{\gamma}\}_{\gamma<\lambda}$, then we set $x^{\lambda}:=y$. Then $x^{\lambda}$ verifies property Pλ and (bλ+1) is satisfied.
    • Assume that $y=x^{\xi_{0}}$ for some $\xi_{0}<\lambda$. By property (bλ) and since $y$ verifies property Pλ, we have that the set $M_{\xi_{0},\lambda}$ is invariant and internally $\mathcal{SC}_{d}$-transitive. Since this set cannot be closed, there is some $z_{1}\in(M_{\xi_{0},\lambda})^{\prime}\setminus M_{\xi_{0},\lambda}$, so that $z_{1}\notin\{x^{\gamma}\}_{\xi_{0}\leq\gamma<\lambda}$. Since $M_{\xi_{0},\lambda}$ is internally $\mathcal{SC}_{d}$-transitive and $z_{1}\in(M_{\xi_{0},\lambda})^{\prime}$, by the triangle inequality it follows that for every $x\in M_{\xi_{0},\lambda}$ there exists a strong $(\epsilon,d)$-chain from $x$ to $z_{1}$ whose points belong to $M_{\xi_{0},\lambda}\cup\{z_{1}\}$. By property (bλ) and using the transitivity of $\mathcal{SC}_{d}$, we have that the point $z_{1}$ verifies property Pλ.
      If $z_{1}=x^{\xi_{1}}$ for some $\xi_{1}<\xi_{0}$, we repeat the same argument replacing $\xi_{1}$ with $\xi_{0}$ and observing that the set $M_{\xi_{1},\lambda}$ is invariant and internally $\mathcal{SC}_{d}$-transitive.
      Since there is no infinitely decreasing sequence of ordinals, we can repeat the previous construction up to a certain $k\in\mathbb{N}$, after which we must have $z_{k+1}\notin\{x^{\gamma}\}_{\gamma<\lambda}$. Then we set $x^{\lambda}:=z_{k+1}$ and since it verifies property Pλ, property (bλ+1) is satisfied.
[…]
  • Case 2.
    $\lambda$ is a limit ordinal.
    Since $|\{x^{\xi}\ |\ \gamma\leq\xi<\lambda\}|\geq\aleph_{0}$ for every $\gamma<\lambda$, by compactness (see [munkres2013topology, Th. 26.9, p. 169]) we can pick
    $y\in\bigcap_{\gamma<\lambda}(\{x^{\xi}\ |\ \gamma\leq\xi<\lambda\})^{\prime}.$
    The point $y$ verifies property Pλ. Indeed, fix $\eta<\lambda$ and $\epsilon>0$ and take $x\in S_{\alpha_{\eta}}(x^{\eta})$. Then there exists $\eta<\xi<\lambda$ such that $d(x^{\xi},y)<\frac{\epsilon}{2}$. By property (bλ), there exists a strong $(\frac{\epsilon}{2},d)$-chain from $x$ to $x^{\xi}$ whose points belong to $M_{\eta,\xi}\cup\{x^{\xi}\}$. By the triangle inequality, it follows that there exists a strong $(\epsilon,d)$-chain from $x$ to $y$ whose points belong to $M_{\eta,\xi}\cup\{y\}\subseteq M_{\eta,\lambda}\cup\{y\}$.
    Consider the following two cases.
    • If $y\notin\{x^{\gamma}\}_{\gamma<\lambda}$, then we set $x^{\lambda}:=y$. Then, $x^{\lambda}$ verifies property Pλ and (bλ+1) is satisfied.
    • Assume that $y=x^{\xi_{0}}$ for some $\xi_{0}<\lambda$. By property (bλ) and since $y$ verifies property Pλ we have that the set $M_{\xi_{0},\lambda}$ is invariant and internally $\mathcal{SC}_{d}$-transitive. Since this set cannot be closed, there exists $z_{1}$ such that $z_{1}\in\left(M_{\xi_{0},\lambda}\right)^{\prime}\setminus M_{\xi_{0},\lambda}.$ Therefore, we have that $z_{1}\notin\{x^{\gamma}\}_{\xi_{0}\leq\gamma<\lambda}$. Since $M_{\xi_{0},\lambda}$ is internally $\mathcal{SC}_{d}$-transitive and $z\in(M_{\xi_{0},\lambda})^{\prime}$, by property (bλ) and by the triangle inequality it follows that for every $x\in M_{\xi_{0},\lambda}$ there exists a strong $(\epsilon,d)$-chain from $x$ to $z_{1}$ whose points belong to $M_{\xi_{0},\lambda}\cup\{z_{1}\}$. By property (bλ) and using the transitivity of $\mathcal{SC}_{d}$, we have that the point $z_{1}$ verifies property Pλ.
      If $z_{1}=x^{\xi_{1}}$ for some $\xi_{1}<\xi_{0}$, we repeat the same argument replacing $\xi_{1}$ with $\xi_{0}$ and observing that the set $M_{\xi_{1},\lambda}$ is invariant and internally $\mathcal{SC}_{d}$-transitive.
      Since there is no infinitely decreasing sequence of ordinals, we can repeat the previous construction up to a certain $k\in\mathbb{N}$, after which we must have $z_{k+1}\notin\{x^{\gamma}\}_{\gamma<\lambda}$. Then we set $x^{\lambda}:=z_{k+1}$ and since it verifies property Pλ, property (bλ+1) is satisfied.
[…]
Assuming that $(X,f)$ does not have a $\mathcal{SC}_{d}$-transitive subsystem, we find that the application
$\gamma\mapsto x^{\gamma}\quad(0\leq\gamma<\lambda)$
is a bijection between $\lambda$ and the set $\{x^{\gamma}\}_{0\leq\gamma<\lambda}$, so we have that $|\{x^{\gamma}\}_{0\leq\gamma<\lambda}|=|\lambda|$. By Hartogs’ Lemma ([hartogs1915problem]), we can take $\lambda$ so large that $|\{x^{\gamma}\}_{0\leq\gamma<\lambda}|>|X|$, which is a contradiction.
Therefore, there exists an ordinal $\nu<\lambda$ such that it is impossible to define the point $x^{\nu}$. More precisely, we have that, if $\nu$ is a successor ordinal, one cannot define the point $x^{\nu}$ if one of the following cases verifies:
  • The set $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is closed, so that $(S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1}))^{\prime}\subseteq S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1}).$ This means that $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is an $\mathcal{SC}_{d}$-transitive subsystem.
  • The set $S_{\beta_{\nu-1},\alpha_{\nu-1}}(x^{\nu-1})$ is not closed and there exists an ordinal $\xi<\nu$ such that the set $M_{\xi,\nu}$ is a closed, invariant and internally $\mathcal{SC}_{d}$-transitive subset, that is a $\mathcal{SC}_{d}$-transitive subsystem.
On the other hand, if $\nu$ is a limit ordinal, then we cannot define the point $x^{\nu}$ if for every
$y\in\bigcap_{\gamma<\nu}(\{x^{\xi}\ |\ \gamma\leq\xi<\nu\})^{\prime},$
we have that $y\in\{x^{\gamma}\}_{\gamma<\nu}$. This implies that there exists an ordinal $\xi<\nu$ such that $M_{\xi,\nu}$ is a closed, invariant and internally $\mathcal{SC}_{d}$-transitive subset, that is a $\mathcal{SC}_{d}$-transitive subsystem. ∎